# Stable Collisionless Tori Around Kerr Black Holes

### v1

@@ Abstract
In low luminosity active galactic nuclei like M87 and Sgr A, the accretion flow in the vicinity of the black hole is in the collisionless regime, meaning that the collisional mean free path of charged particles is much larger than the dynamic length scales. To properly model the particle energization and emission from the collisionless accretion flow, a promising approach is to employ the global general relativistic particle-in-cell simulations—a newly developed, fully kinetic, first-principles method. However, it has been challenging to set up an initial condition that involves collisionless gas with finite angular momentum. We present, for the first time, a class of analytic kinetic equilibria of collisionless tori around a Kerr black hole. We have successfully implemented the collisionless tori in our GPU-based GRPIC code framework Aperture, and found them to be stable for hundreds to thousands of dynamical times in 2D axisymmetric simulations when there is no initial seed magnetic field. These kinetic equilibria serve as ideal starting points for future studies of the physics of collisionless accretion and jet launching.
@@ 1 Introduction
Accretion onto supermassive black holes powers the fascinating phenomena of active galactic nuclei (AGN). As the gas accretes toward the black hole, its gravitational potential energy is converted into heat and radiation, making the nucleus of the galaxy shine in a wide range of electromagnetic spectrum. Some of the accreting black holes can also launch powerful, relativistic jets extending well beyond the galaxy itself, which, in turn, shape the environment of the galaxy.
@@
Recently, the Event Horizon Telescope (EHT) has made unprecedentedly high resolution observations of nearby supermassive black holes, allowing us to directly see the central engines in action. In particular, the EHT obtained polarized images of M87 and Sgr A (Event Horizon Telescope Collaboration et al., 2021a, b, 2024a, 2024b), showing that the bright ring surrounding the black hole shadow is the radiation from the plasma accreting toward the black hole, and that there is a strong magnetic field near the horizon, which may be the key for jet launching. Currently the best model for the horizon-scale emission comes from general relativistic magnetohydrodynamics (GRMHD) simulations (Event Horizon Telescope Collaboration et al., 2019a, 2022). However, GRMHD cannot self-consistently capture the energization of the emitting electrons, leading to significant model uncertainties. Additionally, no model in the GRMHD library was found to satisfactorily reproduce all the observed features of Sgr A. All models tend to produce too much variability (Event Horizon Telescope Collaboration et al., 2022), suggesting that they may be missing some important physical ingredients.
@@
M87 and Sgr A belong to a class of AGN with low luminosity. In these systems, the gas density near the black hole is so low that the plasma is collisionless—the mean free path of Coulomb collisions between charged particles is larger than the system size. Take M87 as an example, a simple one-zone model based on EHT observations gives an estimate of the electron density in the emitting region $n_{e}\sim 10^{4-7}\,{\rm cm}^{-3}$ and electron temperature $T_{e}\sim(1-12)\times 10^{10}\,{\rm K}$ (Event Horizon Telescope Collaboration et al., 2021b). We can then estimate the mean free path for electrons colliding with ions as $\lambda_{\rm mfp}\sim(kT_{e})^{2}/(\pi n_{e}e^{4})\sim 10^{19}T_{e,10}^{2}n_{e,6}^{-1}\,{\rm cm}$ (we use the notation $T_{e,10}\equiv T_{e}/(10^{10}\,{\rm K})$, $n_{e,6}\equiv n_{e}/(10^{6}\,{\rm cm}^{-3})$, same below). Clearly, $\lambda_{\rm mfp}$ is much larger than the gravitational radius $r_{g}\equiv GM/c^{2}\approx 9.6\times 10^{14}\,{\rm cm}$, where we have used a mass of $M=6.5\times 10^{9}M_{\odot}$ for M87 (Event Horizon Telescope Collaboration et al., 2019b). Therefore, the plasma is collisionless. MHD typically assumes that the plasma behaves like an ideal gas that locally reaches thermal equilibrium, but this requires efficient collisions. In the collisionless regime, plasmas do not necessarily maintain local thermal equilibrium; electrons and ions can have quite different distributions. Kinetic effects are important to determine the thermodynamics of the plasma, especially differential heating of electrons and ions, and non-thermal particle acceleration, which cannot be captured in MHD simulations.
@@
Recently, significant progress has been made in developing a first-principles, kinetic simulation method—general relativistic particle-in-cell (GRPIC)—to study the collisionless plasma dynamics around black holes (e.g., K. Parfrey et al., 2019; B. Crinquand et al., 2020, 2021, 2022; A. Bransgrove et al., 2021; I. El Mellah et al., 2022, 2023; J. Mehlhaff et al., 2025; A. Y. Chen et al., 2025; Y. Yuan et al., 2025). The GRPIC approach follows the evolution of the electromagnetic field and particle distribution function self-consistently by solving the coupled Maxwell and Vlasov equations in the curved spacetime. The approach retains the full information of collective plasma behavior as well as nonthermal particle acceleration. Global GRPIC simulations have just started to be employed to study the collisionless accretion process (A. Galishnikova et al., 2023; J. Vos et al., 2025), which demonstrate different behaviors compared to MHD, particularly with respect to magnetic reconnection, instabilities caused by plasma anisotropy, heat flux, and plasma mixing at the jet/disk boundary. However, the studies so far have only considered accretion of zero-angular-momentum gas, which is an over simplification—in reality, the gas reservoir at the center of galaxies likely has nonzero angular momentum. To understand the accretion physics in low luminosity AGN, we need to focus on more realistic configurations.
@@
We would like to build a self-consistent model of collisionless accretion flows, starting from a physical initial condition where the gas supply near the black hole has finite angular momentum. A commonly used initial condition in GRMHD simulations is the the Fishbone-Moncrief torus (L. G. Fishbone & V. Moncrief, 1976), which is a hydrodynamic equilibrium when there is no magnetic field. However, initializing such a torus in the collisionless regime already presents as the first challenge. For the Fishbone-Moncrief torus, the equilibrium is possible because a combination of the gravitational force and the pressure gradient provides the centripetal force for the orbiting gas element. The pressure, temperature and enthalpy of the gas is specified based on the adiabatic ideal gas law, assuming local thermal equilibrium. In the collisionless regime, if we set up a plasma distribution that is locally Maxwellian following the temperature profile of the Fishbone-Moncrief torus, it turns out that the distribution quickly evolves away from the initial configuration, even when electromagnetism is turned off. This is because particles are freely streaming along geodesics and there is no way to maintain a local thermal equilibrium which is required by an adiabatic ideal gas.
@@
In this Letter, we present, for the first time, a class of analytic kinetic equilibria of collisionless tori around a Kerr black hole. We have successfully implemented the collisionless tori in our GPU-based GRPIC code framework Aperture, and found them to be stable for hundreds to thousands of dynamical times in 2D axisymmetric simulations when there is no initial seed magnetic field. These kinetic equilibria serve as ideal testbeds for us to further study the physics of collisionless accretion and jet launching.
@@
Figure 1: Cross sections of the density distribution for three axisymmetric tori. Panels I, II, and III correspond to the initial tori with inner radii $r_{\mathrm{in}}=11r_{g}$, $10r_{g}$, and $9r_{g}$, respectively, each normalized to unity. Panels IV, V, and VI are these same tori after evolving for $800r_{g}$. We include electromagnetic effects but there is no initial background field. All cases have temperature $T=10^{-3}m$ and constant angular momentum $L_{0}\approx 3.986m$, corresponding to $r_{0}=14r_{g}$ around a black hole with spin $a=0.999$. The fluctuations seen in the right column are consistent with particle shot-noise for 20 particles per cell. In each panel, the black circle shows the black hole event horizon, $r_{h}=r_{g}(1+\sqrt{1-a^{2}})$.
@@ 2.1 Vlasov Equation and Jeans’s Theorem
To describe a collisionless ensemble of particles in general relativity, we use a distribution function $f(x^{\mu},p_{\mu})$ which gives the density of particles in phase space (F. A. Rasio et al., 1989). Here $x^{\mu}=(x^{0},x^{1},x^{2},x^{3})$ is the position 4-vector, $x^{0}=t$ is the coordinate time, and $p_{\mu}=(p_{0},p_{1},p_{2},p_{3})$ is the covariant 4-momentum. We use geometrized units such that $c=G=k_{B}=1$; Greek indices take on values $0,1,2,3$ while Latin indices take on values $1,2,3$. Specifically, the number of particles $dN$ in a phase space volume element $dV_{x}dV_{p}$ is given by
@@
This is the collisionless Boltzmann equation, or the Vlasov equation. In coordinate form, Equation (4) can be written as
@@
In Kerr spacetime, test particle trajectories have four constants of motion. With the symmetries of the Kerr metric, Hamilton’s equations (3) provide two constants of motion, energy $E=-p_{0}\$and angular momentum $L_{z}=p_{\phi}$. The third constant of motion is the particle’s rest mass $m=(-p^{\mu}p_{\mu})^{1/2}$. B. Carter (1968) showed that there is a fourth constant $Q$ now called Carter’s constant derived in Appendix C. This allows us to write the Boyer-Lindquist momentum $p_{\mu}$ as a function of these constants and $r,\theta$
@@ 2.2 Collisionless Torus Distribution
To write down the analytic solutions, we work in the Boyer-Lindquist coordinates and use the $3+1$ formalism  (S. S. Komissarov, 2004, and Appendix B). It is well known that at constant angular momentum there is a range of energies that correspond to stable geodesic orbits (J. Levin & G. Perez-Giz, 2008). We choose a distribution function with constant angular momentum $L_{0}$ and maximum energy $E_{\max}$ that confines the geodesics to a finite region. We choose a Maxwell-Boltzmann distribution in $E$ with exponential scale $T$  (as in P. Rioseco & O. Sarbach, 2017; P. Li et al., 2023)
@@ 3.1 Change of Variables
To facilitate integration over momentum space and enable efficient sampling of particle velocities, we introduce polar coordinates in momentum space (in Boyer-Lindquist coordinates):
@@ 4 Simulations
We use our GPU-based GRPIC code framework Aperture [^1] The full code, including the implementation of the torus setup introduced in this paper, is hosted at https://github.com/fizban007/Aperture4. (A. Y. Chen et al., 2025) to follow the evolution of the analytic, collisionless torus initial condition. For simplicity, we consider an electron-positron pair plasma. The simulations are carried out in 2D axisymmetry using Kerr-Schild spherical coordinates, adopting the $3+1$ formalism of S. S. Komissarov (2004). The benefit of this coordinate system is that there is no singularity at the event horizon, unlike the Boyer-Lindquist coordinates, and we can choose the domain boundary to be within the horizon. Such is also the choice of other GRPIC codes such as K. Parfrey et al. (2019); C. Meringolo et al. (2025). However, the formalism developed in Section 3 is based on the Boyer-Lindquist coordinates, therefore a coordinate transformation is required to convert the relevant quantities to their corresponding Kerr-Schild versions. For the density, $S^{0}$ turns out to be the same in Boyer-Lindquist and Kerr-Schild coordinates due to Equation (A2), therefore we can obtain the mass and charge densities in Kerr-Schild coordinates as
@@
When initializing particles in the simulations, we inject a fixed number of particles per cell (ppc) in the torus where $\rho_{m,s}^{\rm(KS)}>0$, and assign each particle a weight proportional to the mass density $\rho_{m,s}^{\rm(KS)}$. To ensure that the initial charge and current density is zero, we inject electrons and positrons in pairs with identical location and momenta. For particle momenta, we first sample the Boyer-Lindquist momentum $p_{\mu}^{\rm(BL)}$ according to §3.3, then transform to Kerr-Schild coordinates following Equation (A3). The left column of Figure 1 shows the initial total mass density for a few different realizations of the collisionless torus. It can be seen that with fixed $r_{0}$ or $L_{0}$, a decreasing $r_{\rm in}$, or equivalently, an increasing $E_{\max}$, leads to a larger torus. Note that the density peak is not necessarily located at $r_{0}$; it is usually at $r_{\rm peak}<r_{0}$.
@@
We perform a series of tests for the equilibrium torus. The simulations do not include any initial electric or magnetic field, but we do evolve all 6 components of the electromagnetic field. The computational domain extends from just inside the event horizon to $r\approx 60\,r_{g}$, well beyond the outer boundary of the torus. The grid is logarithmically spaced in the radial direction and uniform in the $\theta$ direction. We use a numerical resolution of $N_{r}\times N_{\theta}=1024\times 1024$ and initialize 20 particles per cell. To properly capture the kinetic physics, we need to resolve the plasma skin depth $\lambda_{p}=c/\omega_{p}=\sqrt{m^{2}c^{2}/(4\pi\rho_{m}e^{2})}$ and the Debye length $\lambda_{D}=\sqrt{k_{B}T_{\rm eff}m/(4\pi\rho_{m}e^{2})}$, where $T_{\rm eff}$ is the effective temperature of the plasma in its rest frame, to be defined in Section 5. Since our effective temperature is typically non-relativistic, $k_{B}T_{\rm eff}/mc^{2}\ll 1$, the smallest scale we need to resolve is the Debye length. If the Debye length were not resolved, the plasma would be numerically heated up until the Debye length gets above the grid scale, modifying the initial distribution and generally inflating the torus. In our simulations, we have chosen a density such that at the density peak, the Debye length is $0.1r_{g}$ and the plasma skin depth is $(10)^{1/2}r_{g}$, well resolved by our grid. Toward the edge of the torus, both the effective temperature and the density drops to zero, therefore it is not immediately clear whether $\lambda_{D}$ grows closer to the edge. In practice, we find that resolving the Debye length at the density peak seems sufficient, and there is no significant numerical heating over a long simulation. We evolve the system for several hundred dynamical times. The final densities for the three cases are shown in the right column of Figure 1. There is no noticeable evolution in the bulk of the torus; the appearance of noise near the edge of the torus is due to the low density and relatively high particle shot-noise. In particular, there is no inflation of the torus, confirming that no significant numerical heating has occurred. We thus conclude that the initial condition is indeed a steady state.

### v2

@@ Abstract
In low-luminosity active galactic nuclei like M87 and Sgr A, the accretion flow in the vicinity of the black hole is in the collisionless regime, meaning that the collisional mean free path of charged particles is much larger than the dynamical length scales. To properly model the particle energization and emission from the collisionless accretion flow, a promising approach is to employ the global general-relativistic particle-in-cell simulations—a newly developed, fully kinetic, first-principles method. However, it has been challenging to set up an initial condition that involves collisionless gas with finite angular momentum. We present, for the first time, a class of analytic kinetic equilibria of collisionless tori around a Kerr black hole. We have successfully implemented the collisionless tori in our GPU-based GRPIC code framework Aperture, and found them to be stable for hundreds to thousands of dynamical times in 2D axisymmetric simulations when there is no initial seed magnetic field. These kinetic equilibria serve as ideal starting points for future studies of the physics of collisionless accretion and jet launching.
@@ 1 Introduction
Accretion onto supermassive black holes powers the fascinating phenomena of active galactic nuclei (AGN). As the gas accretes toward the black hole, its gravitational potential energy is converted into heat and radiation, making the nucleus of the galaxy shine across a wide range of the electromagnetic spectrum. Some of the accreting black holes can also launch powerful, relativistic jets extending well beyond the galaxy itself, which, in turn, shape the environment of the galaxy.
@@
Recently, the Event Horizon Telescope (EHT) has made unprecedented, high-resolution observations of nearby supermassive black holes, allowing us to directly see the central engines in action. In particular, the EHT obtained polarized images of M87 and Sgr A (Event Horizon Telescope Collaboration et al., 2021a, b, 2024a, 2024b), showing that the bright ring surrounding the black hole shadow is the radiation from the plasma accreting toward the black hole, and that there is a strong magnetic field near the horizon, which may be the key for jet launching. Currently, the best models for the horizon-scale emission come from general relativistic magnetohydrodynamics (GRMHD) simulations (Event Horizon Telescope Collaboration et al., 2019a, 2022). However, GRMHD cannot self-consistently capture the energization of the emitting electrons, leading to significant model uncertainties. Additionally, no model in the GRMHD library has been found to satisfactorily reproduce all the observed features of Sgr A. These models cannot simultaneously explain the variability and geometric properties of this source (Event Horizon Telescope Collaboration et al., 2022), suggesting that they may be missing some important physical ingredients.
@@
M87 and Sgr A belong to a class of AGN with low-luminosity. In these systems, the gas density near the black hole is so low that the plasma is collisionless—the mean free path of Coulomb collisions between charged particles is larger than the system size. Take M87 as an example, a simple one-zone model based on EHT observations gives an estimate of the electron density in the emitting region $n_{e}\sim 10^{4\text{--}7}\,{\rm cm}^{-3}$ and electron temperature $T_{e}\sim(1\text{--}12)\times 10^{10}\,{\rm K}$ (Event Horizon Telescope Collaboration et al., 2021b). We can then estimate the mean free path for electrons colliding with ions as $\lambda_{\rm mfp}\sim(kT_{e})^{2}/(\pi n_{e}e^{4})\sim 10^{19}T_{e,10}^{2}n_{e,6}^{-1}\,{\rm cm}$ (we use the notation $T_{e,10}\equiv T_{e}/(10^{10}\,{\rm K})$, $n_{e,6}\equiv n_{e}/(10^{6}\,{\rm cm}^{-3})$). Clearly, $\lambda_{\rm mfp}$ is much larger than the gravitational radius $r_{g}\equiv GM/c^{2}\approx 9.6\times 10^{14}\,{\rm cm}$, where we have used a mass of $M=6.5\times 10^{9}M_{\odot}$ for M87 (Event Horizon Telescope Collaboration et al., 2019b). Therefore, the plasma is collisionless. MHD typically assumes that the plasma behaves like an ideal gas that locally reaches thermal equilibrium, but this requires efficient collisions. In the collisionless regime, plasmas do not necessarily maintain local thermal equilibrium; electrons and ions can have quite different distributions. Kinetic effects are important to determine the thermodynamics of the plasma, especially differential heating of electrons and ions, and non-thermal particle acceleration, which cannot be captured in MHD simulations.
@@
Recently, significant progress has been made in developing a first-principles, kinetic simulation method—general relativistic particle-in-cell (GRPIC)—to study the collisionless plasma dynamics around black holes (e.g., K. Parfrey et al., 2019; B. Crinquand et al., 2020, 2021, 2022; A. Bransgrove et al., 2021; I. El Mellah et al., 2022, 2023; J. Mehlhaff et al., 2025; A. Y. Chen et al., 2025; Y. Yuan et al., 2025). The GRPIC approach follows the evolution of the electromagnetic field and particle distribution function self-consistently by solving the coupled Maxwell and Vlasov equations in the curved spacetime. The approach retains the full information of collective plasma behavior as well as nonthermal particle acceleration. Global GRPIC simulations have just started to be employed to study the collisionless accretion process (A. Galishnikova et al., 2023; J. Vos et al., 2025), which demonstrate different behaviors compared to MHD, particularly with respect to magnetic reconnection, instabilities caused by plasma anisotropy, heat flux, and plasma mixing at the jet/disk boundary. However, the studies so far have only considered accretion of zero-angular-momentum gas, which is an oversimplification—in reality, the gas reservoir at the center of galaxies likely has nonzero angular momentum. To understand the accretion physics in low-luminosity AGN, we need to focus on more realistic configurations.
@@
We would like to build a self-consistent model of collisionless accretion flows, starting from a physical initial condition where the gas supply near the black hole has finite angular momentum. A commonly used initial condition in GRMHD simulations is a hydrodynamic equilibrium assuming a perfect fluid (e.g., L. G. Fishbone & V. Moncrief, 1976; M. Abramowicz et al., 1978). However, initializing such a torus in the collisionless regime already presents the first challenge. For the perfect fluid torus, the equilibrium is possible because a combination of the gravitational force and the pressure gradient provide the centripetal force for the orbiting gas element. The perfect fluid assumption requires that the stress-energy tensor takes the simple form $\mathrm{diag}(\varepsilon,P,P,P)$ in the fluid rest frame, where $\varepsilon$ is the total energy density and $P$ is the pressure.
In the collisionless regime, if we set up a plasma distribution that is locally Maxwellian following an instance of the Fishbone-Moncrief torus (e.g., C. F. Gammie et al., 2003), it turns out that the distribution quickly evolves away from the initial configuration, even when electromagnetism is turned off. This is because particles are freely streaming along geodesics and there is no way to maintain the perfect fluid assumption as required by these hydrodynamic torus solutions.
@@
In this paper, we present, for the first time, a class of analytic kinetic equilibria of collisionless tori around a Kerr black hole. We have successfully implemented the collisionless tori in our GPU-based GRPIC code framework Aperture, and found them to be stable for hundreds to thousands of dynamical times in 2D axisymmetric simulations when there is no initial seed magnetic field. These kinetic equilibria serve as ideal testbeds for us to further study the physics of collisionless accretion and jet launching.
@@
Figure 1: Cross sections of the density distribution for three axisymmetric tori. Panels I, II, and III correspond to the initial tori with inner radii $r_{\mathrm{in}}=11\,r_{g}$, $10\,r_{g}$, and $9\,r_{g}$, respectively, each normalized to unity. Panels IV, V, and VI show these same tori after evolving for $800\,r_{g}/c$. We include electromagnetic effects but there is no initial background field. All cases have temperature $T=10^{-3}m$ and constant angular momentum $L_{0}\approx 3.986\,m$, corresponding to $r_{0}=14\,r_{g}$ around a black hole with spin $a=0.999$. The fluctuations seen in the right column are consistent with particle shot noise for 20 particles per cell. In each panel, the black circle shows the black hole event horizon, $r_{h}=r_{g}(1+\sqrt{1-a^{2}})$.
@@ 2.1 Vlasov Equation and Jeans’s Theorem
To describe a collisionless ensemble of particles in general relativity, we use a distribution function $f(x^{\mu},p_{\mu})$ which gives the density of particles in phase space (F. A. Rasio et al., 1989). Here $x^{\mu}=(x^{0},x^{1},x^{2},x^{3})$ is the position 4-vector, $x^{0}=t$ is the coordinate time, and $p_{\mu}=(p_{0},p_{1},p_{2},p_{3})$ is the covariant 4-momentum. We use geometrized units such that $c=G=k_{B}=1$; Greek indices take on values $0,1,2,3$ while Latin indices take on values $1,2,3$. Throughout this paper, we use three different coordinate systems: Boyer–Lindquist (BL), Kerr–Schild (KS) and fluid rest frame (FRF). BL coordinates are most convenient for the analytic calculations, KS coordinates are used in the simulation because there is no coordinate singularity at the horizon and the FRF is required when calculating any local properties of the gas. Now, the number of particles $dN$ in a phase space volume element $dV_{x}dV_{p}$ is given by
@@
This is the collisionless Boltzmann (Vlasov) equation. In coordinate form, Equation (4) can be written as
@@
In Kerr spacetime, test particle trajectories have four constants of motion. With the symmetries of the Kerr metric, Hamilton’s equations (3) provide two constants of motion, energy $E=-p_{0}\$and angular momentum $L_{z}=p_{\phi}$. The third constant of motion is the particle’s rest mass $m=(-p^{\mu}p_{\mu})^{1/2}$. B. Carter (1968) showed that there is a fourth constant $Q$ now called Carter’s constant derived in Appendix C. This allows us to write the Boyer–Lindquist momentum $p_{\mu}$ as a function of these constants and $r,\theta$
@@ 2.2 Collisionless Torus Distribution
To write down the analytic solutions, we work in the Boyer–Lindquist coordinates and use the $3+1$ formalism  (S. S. Komissarov, 2004, and Appendix B). It is well known that, at constant angular momentum, there is a range of energies that correspond to stable geodesic orbits  (e.g., J. M. Bardeen et al., 1972; J. Levin & G. Perez-Giz, 2008). We choose a distribution function with constant angular momentum $L_{0}$ and maximum energy $E_{\max}$ that confines the geodesics to a finite region. We choose a Maxwell-Boltzmann distribution in $E$ with exponential scale $T$  (as in P. Rioseco & O. Sarbach, 2017; P. Li et al., 2023)
@@ 3.1 Change of Variables
To facilitate integration over momentum space and enable efficient sampling of particle velocities, we introduce polar coordinates in momentum space (in Boyer–Lindquist coordinates):
@@ 4 Simulations
We use our GPU-based GRPIC code framework Aperture [^1] The full code, including the implementation of the torus setup introduced in this paper, is hosted at https://github.com/fizban007/Aperture4. (A. Y. Chen et al., 2025) to follow the evolution of the analytic, collisionless torus initial condition. For simplicity, we consider an electron-positron pair plasma. The simulations are carried out in 2D axisymmetry using Kerr–Schild spherical coordinates, adopting the $3+1$ formalism of S. S. Komissarov (2004). The benefit of this coordinate system is that there is no singularity at the event horizon, unlike the Boyer–Lindquist coordinates, and we can choose the domain boundary to be within the horizon. Similar choices are made in other GRPIC codes such as K. Parfrey et al. (2019); C. Meringolo et al. (2025). However, the formalism developed in Section 3 is based on the Boyer–Lindquist coordinates, therefore a coordinate transformation is required to convert the relevant quantities to their corresponding Kerr–Schild versions. For the density, $S^{0}$ turns out to be the same in Boyer–Lindquist and Kerr–Schild coordinates due to Equation (A2), therefore we can obtain the mass and charge densities in Kerr–Schild coordinates as
@@
When initializing particles in the simulations, we inject a fixed number of particles per cell (ppc) in the torus where $\rho_{m,s}^{\rm(KS)}>0$, and assign each particle a weight proportional to the mass density $\rho_{m,s}^{\rm(KS)}$. To ensure that the initial charge and current density is zero, we inject electrons and positrons in pairs with identical location and momenta. For particle momenta, we first sample the Boyer–Lindquist momentum $p_{\mu}^{\rm(BL)}$ according to Section 3.3, then transform to Kerr–Schild coordinates following Equation (A3). The left column of Figure 1 shows the initial total mass density for a few different realizations of the collisionless torus. It can be seen that with fixed $r_{0}$ or $L_{0}$, a decreasing $r_{\rm in}$, or equivalently, an increasing $E_{\max}$, leads to a larger torus. Note that the density peak is not necessarily located at $r_{0}$; it is usually at $r_{\rm peak}<r_{0}$.
@@
We perform a series of tests for the equilibrium torus. The simulations do not include any initial electric or magnetic field, but we do evolve all 6 components of the electromagnetic field. The computational domain extends from just inside the event horizon to $r\approx 60\,r_{g}$, well beyond the outer boundary of the torus. The grid is logarithmically spaced in the radial direction and uniform in the $\theta$ direction. We use a numerical resolution of $N_{r}\times N_{\theta}=1024\times 1024$ and initialize 20 particles per cell. To properly capture the kinetic physics, we need to resolve the plasma skin depth $\lambda_{p}=c/\omega_{p}=\sqrt{m^{2}c^{2}/(4\pi\rho_{m}e^{2})}$ and the Debye length $\lambda_{D}=\sqrt{k_{B}T_{\rm eff}m/(4\pi\rho_{m}e^{2})}$, where $T_{\rm eff}$ is the effective temperature of the plasma in its rest frame, to be defined in Section 5. Since our effective temperature is typically non-relativistic, $k_{B}T_{\rm eff}/mc^{2}\ll 1$, the smallest scale we need to resolve is the Debye length. If the Debye length were not resolved, the plasma would be numerically heated up until the Debye length gets above the grid scale, modifying the initial distribution and generally inflating the torus. In our simulations, we have chosen a density such that at the density peak, the Debye length is $0.1r_{g}$ and the plasma skin depth is $\sqrt{10}\,r_{g}$, well resolved by our grid. This resolution is sufficient to resolve the relevant length scales throughout the torus, despite the fact that the position where the Debye length is least resolved does not coincide exactly with the density peak, and depends on the radial extent of our grid.
We evolve the system for several hundred dynamical times. The final densities for the three cases are shown in the right column of Figure 1. There is no noticeable evolution in the bulk of the torus; the appearance of noise near the edge of the torus is due to the low density and relatively high particle shot-noise. In particular, there is no inflation of the torus, confirming that no significant numerical heating has occurred. We thus conclude that the initial condition is indeed a steady state.